Jorge pushes a 10 kg box of medicine along a rough surface with 35N of force. He observes that it is only accelerating at a rate of 2 m/s2. What must the value of the force of friction be? (Assume the only two forces acting on the object are friction and Jorge).

 



1
Expert's answer
2021-03-01T12:39:08-0500

"F-F_{fr}=ma, \implies"

"F_{fr}=F-ma,"

"F_{fr}=35-10\cdot 2=15~\text{N}."